OpenStudy (anonymous):

A poker hand consists of 5 cards dealt at random without replacement from a standard deck of 52 cards of which 26 are red and the rest black. A poker hand is dealt. Find the chance that the hand contains three red cards and two black cards

OpenStudy (anonymous):

number of ways to get 3 red cards = 2600 number of ways to get 2 black cards = 325 Total number of ways you could get 3 red and 2 black = 845,000 total number of possiblities in 5 card poker = 2,598,960 so answer = 845000/2598960 0.325130052

OpenStudy (anonymous):

great could break down this for me in fractions please so that i compare with what i had previously?

OpenStudy (anonymous):

number of ways to get 3 red cards \[\left(\begin{matrix}26 \\3\end{matrix}\right)\]= 2600

OpenStudy (anonymous):

number of ways to get 2 black cards \[\left(\begin{matrix}26 \\ 2\end{matrix}\right)\]= 325

OpenStudy (anonymous):

Total number of ways you could get 3 red and 2 black =2600*325=845,000

OpenStudy (anonymous):

total number of possiblities in 5 card poker\[\left(\begin{matrix}52 \\ 5\end{matrix}\right)\] = 2,598,960

OpenStudy (anonymous):

so answer = 845000/2598960 =0.325130052

OpenStudy (anonymous):

I was stuck with this question, thank you!